40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Feb 4 20:13:39 2019, gautam, Update, SUS, ETMY suspension oddness ETMY_sensors_1_Feb_2019_2230_PST.pdfETMY_UL.pdfETMY_sensors_timeDomain.pdf
    Reply  Thu Feb 7 19:34:18 2019, gautam, Update, SUS, ETMY suspension oddness ETMY_sensors.pdfpostDiag.pdf
       Reply  Fri Feb 8 20:48:52 2019, gautam, Update, LSC, IFO recovery IMG_7307.JPG
          Reply  Mon Feb 11 15:41:49 2019, gautam, Update, LSC, TRY 60 Hz solved - but clipping persists 
             Reply  Mon Feb 11 16:38:34 2019, gautam, Update, LSC, ETMY OL calibration updated OL_calib_ETMY_PERROR.pdfOL_calib_ETMY_YERROR.pdf
             Reply  Fri Feb 15 18:42:57 2019, gautam, Update, LSC, TRY 60 Hz solved  IMG_7323.JPG
Message ID: 14459     Entry time: Fri Feb 15 18:42:57 2019     In reply to: 14446
Author: gautam 
Type: Update 
Category: LSC 
Subject: TRY 60 Hz solved  

A more permanent fix than a crocodile clip was implemented. Should probably look to do this for the X end unit as well.

Attachment 1: IMG_7323.JPG  1.892 MB  | Hide | Hide all
IMG_7323.JPG
ELOG V3.1.3-